# Dialogues with AI Reduce Beliefs in Misinformation but Build No Lasting Discernment Skills

### v1

@@ 8.1. Categories and Examples
Our analysis identified three primary behavioral categories that characterize Human-AI conversational dynamics. These categories could help explain potential mechanisms underlying the performance degradation of misinformation discernment accuracy over time: increased agreement with the system over time, reduced questioning behavior, and decreased independent thinking. Agreement behavior captures instances where one party validates the other’s perspective, distinguishing between human agreement with AI versus AI agreement with human responses. People typically use casual affirmations (“you’re right,” “that makes sense,” “totally”) while AI employs more formal validation (“you make a valid point,” “your perspective is valid,” “I agree with your assessment”). Independent thinking encompasses expressions of personal reasoning and uncertainty. People demonstrate subjective stance-taking (“I think,” “I believe,” “maybe”) compared to AI’s analytical hedging (“it seems,” “this suggests,” “considering”). Probing behavior reflects information-seeking and curiosity. People ask direct, challenging questions (“why do you,” “how do you know,” “are you sure”) while AI engages in Socratic inquiry (“what do you think,” “help me understand,” “have you considered”). We include the complete keyword list in the Appendix section E.3 Figure 16, 17, 18.
@@ 8.2. Insights
Figure 8 shows the longitudinal analysis of conversations, revealing shifts in conversational dynamics across the month-long period. Human agreement with AI increased from 23.0% in Week 0 to 28.5% in Week 4, while AI agreement with humans remained relatively stable (4.2% to 6.4%), suggesting growing human acceptance of AI perspectives.
Conversely, human independent thinking remained consistently low across all weeks (7.2% to 7.6%), while AI independent thinking increased from 19.8% to 25.1%, indicating that the AI maintained analytical autonomy while humans’ independent thinking ability did not vary much. The probing behavior showed a decline in curiosity over time, with human questioning of AI increasing slightly (from 15.6% to 19.9%) and then declining to 17.6% in Week 4, while AI probing of humans remained relatively stable (25.6% to 24.6%).
These conversational patterns align with our broader findings that AI assistance provided immediate performance benefits during AI-assisted sessions, but participants’ independent reasoning abilities deteriorated over the month-long study period. Together, this suggests a potential overreliance on AI guidance at the expense of fake news discernment skills for misinformation detection.
@@ 10. Discussion
Our findings reveal a critical distinction between correcting false beliefs and building lasting detection capabilities. Previous work by Costello et al. (Costello et al., 2024) demonstrated that AI dialogues can reduce belief in conspiracy theories by 22%. Our results confirm that AI can effectively change beliefs about specific misinformation items with participants showing an average of 21% improvement in accuracy during AI-assisted sessions across all weeks. However, this immediate belief correction did not translate into improved independent discernment abilities. While participants benefited from AI’s persuasive reasoning (+21 pp accuracy gain), they later became less able to evaluate new misinformation on their own (–15.3 pp decline by week 4). Interestingly, this decline was concentrated on detecting fake content, suggesting that once participants learned to trust AI for verification, they lost vigilance in scrutinizing suspicious claims themselves. This has implications for how we currently use AI to assist us in information processing. If the goal is simply to reduce the belief in specific falsehoods “in the moment,” AI can be effective. But if the goal is to build people’s skills to recognize falsehoods on their own, our results suggest current AI approaches may fail—or even backfire by fostering reliance.
@@
The temporal dynamics of our findings align with broader research on Human-AI collaboration, showing immediate performance benefits that mask underlying skill degradation. Similar patterns have been observed in other domains where AI assistance provides cognitive scaffolding — users often experience enhanced performance during assisted trials but then develop dependency that undermines their capabilities to think on their own (Danry et al., 2023; Agarwal et al., 2025; Kosmyna et al., 2025). Similarly, a recent study also found that doctors using AI classification tools to detect cancer became worse at detecting cancer alone after having used the tool (Budzyń et al., 2025).
@@ 11. Limitations and Future Work
Study Duration. While our month-long study captured important longitudinal trends, longer-term studies would be valuable to understand whether the observed dependency effects could persist, worsen, or eventually reverse after extended periods without AI assistance.
Several promising avenues emerge from our limitations. Longitudinal studies extending beyond four weeks could reveal whether skill degradation continues or whether recovery occurs after AI assistance withdrawal. Cross-cultural validation would establish whether dependency effects vary across different educational systems, media landscapes, or cultural approaches to authority and expertise. Comparative studies examining different AI pedagogical approaches — particularly Socratic questioning methods that prompt user reasoning rather than providing direct answers — could identify design principles that enhance rather than undermine skill development.
@@ 12. Conclusion
This study reveals a concerning paradox in AI-assisted misinformation detection: while persuasive AI systems produce immediate accuracy gains (+21%), they simultaneously undermine long-term independent judgment (-15.3% by week 4). Our longitudinal analysis demonstrates that current approaches prioritize belief correction over skill development, creating dependency rather than durable discernment capabilities. The findings challenge assumptions about AI’s educational potential and highlight the need for systems designed to enhance rather than replace human reasoning. As AI becomes increasingly sophisticated, ensuring these tools build critical thinking skills rather than cognitive dependency becomes essential for maintaining public resilience to misinformation.
@@ D.2. Conversation Characteristics
The conversations typically followed a pattern where participants would state their initial belief about a news item’s authenticity, the AI would probe their reasoning, provide evidence or counterevidence, and engage in iterative dialogue aimed at improving the participant’s detection accuracy. The relatively consistent conversation length across weeks indicates stable engagement patterns throughout the longitudinal study. Some of the examples of these conversations are present in Figure 12, 13 and 14.

### v2

@@ 8.1. Categories and Examples
Our analysis identified three primary behavioral categories that characterize Human-AI conversational dynamics. These categories could help explain potential mechanisms underlying the performance degradation of misinformation discernment accuracy over time: increased agreement with the system over time, reduced questioning behavior, and decreased independent thinking. Agreement behavior captures instances where one party validates the other’s perspective, distinguishing between human agreement with AI versus AI agreement with human responses. People typically use casual affirmations (“you’re right,” “that makes sense,” “totally”) while AI employs more formal validation (“you make a valid point,” “your perspective is valid,” “I agree with your assessment”). Independent thinking encompasses expressions of personal reasoning and uncertainty. People demonstrate subjective stance-taking (“I think,” “I believe,” “maybe”) compared to AI’s analytical hedging (“it seems,” “this suggests,” “considering”). Probing behavior reflects information-seeking and curiosity. People ask direct, challenging questions (“why do you,” “how do you know,” “are you sure”) while AI engages in Socratic inquiry (“what do you think,” “help me understand,” “have you considered”). We include the complete keyword list in the Appendix section E.3 Figure 17, 18, 19.
@@ 8.2. Insights
Figure 9 shows the longitudinal analysis of conversations, revealing shifts in conversational dynamics across the month-long period. Human agreement with AI increased from 20.9% in Week 0 to 28.5% in Week 4 (Kruskal–Wallis ; Week 0 vs. Week 2 and Week 0 vs. Week 4 both , Week 2 vs. Week 4 ), while AI agreement with humans increased modestly from 3.9% to 6.6% (Kruskal–Wallis ; Week 0 vs. Week 2 and Week 0 vs. Week 4 both , Week 2 vs. Week 4 ), suggesting growing human acceptance of AI perspectives over time alongside a small but reliable increase in reciprocal alignment.
Conversely, human independent thinking remained consistently low and statistically unchanged across all weeks (Week 0: , Week 2: , Week 4: ; Kruskal–Wallis ), while AI independent thinking increased from 16.9% to 24.8% (Week 0: , Week 2: , Week 4: ; Kruskal–Wallis ; Week 0 vs. Week 2 and Week 0 vs. Week 4 both , Week 2 vs. Week 4 ), indicating that the AI maintained and strengthened analytical autonomy while humans’ independent thinking did not measurably change. Probing behavior showed a more nuanced pattern over time: human questioning of AI increased from 12.8% to 17.7% and then declined slightly to 15.6% in Week 4 (Kruskal–Wallis ; only Week 0 vs. Week 2 reached significance, ), while AI probing of humans decreased from 36.0% to around 25% by Week 2 and then remained relatively stable (Week 0: , Week 2: , Week 4: ; Kruskal–Wallis ; Week 0 vs. Week 2 , Week 0 vs. Week 4 , Week 2 vs. Week 4 ).
These conversational patterns align with our broader findings that AI assistance provided immediate performance benefits during AI-assisted sessions, but participants’ independent conversational reasoning, as captured by our independent-thinking and probing metrics, did not seem to improve over the month-long study period, even as agreement with the AI increased. Together with the performance results reported above, this pattern is consistent with a potential overreliance on AI guidance at the expense of developing robust fake news discernment skills for misinformation detection.
@@ 10. Discussion
Our findings reveal a critical distinction between correcting false beliefs and building lasting detection capabilities. Previous work by Costello et al. (Costello et al., 2024) demonstrated that AI dialogues can reduce belief in conspiracy theories by 22%. Our results confirm that AI can effectively change beliefs about specific misinformation items with participants showing an average of 21% improvement in accuracy during AI-assisted sessions across all weeks. However, this immediate belief correction did not translate into improved independent discernment abilities. While participants benefited from AI’s persuasive reasoning (+21 pp accuracy gain), they later became less able to evaluate new misinformation on their own immediately after using the AI (–15.3 pp decline by week 4) and had no significant change in their independent abilities before interacting with AI. Interestingly, the decline after interacting with the AI was concentrated on detecting fake content, suggesting that once participants learned to trust AI for verification, they may have lost vigilance in scrutinizing suspicious claims themselves. This has implications for how we currently use AI to assist us in information processing. If the goal is simply to reduce the belief in specific falsehoods “in the moment,” AI can be effective. But if the goal is to build people’s skills to recognize falsehoods on their own, our results suggest current AI approaches may fail—or even backfire by fostering reliance.
@@
The temporal dynamics of our findings align with broader research on Human-AI collaboration, showing immediate performance benefits that may mask underlying skill degradation or reliance on AI. Similar patterns have been observed in other domains where AI assistance provides cognitive offloading — users often experience enhanced performance during AI-assisted trials but then develop dependency that undermines their capabilities to think or act on their own (Danry et al., 2023; Agarwal et al., 2025; Kosmyna et al., 2025). Similarly, a recent study also found that doctors using AI classification tools to detect cancer became worse at detecting cancer alone after having used the tool (Budzyń et al., 2025).
@@ 11. Limitations and Future Work
Study Duration. Longitudinal studies extending beyond four weeks could reveal whether the observed reliance effect persists, whether independent detection abilities recover after extended AI withdrawal, or whether skills continue to deteriorate. Research on cognitive offloading in other domains shows mixed outcomes: while some skills recover after tool removal (Grinschgl et al., 2021), others show persistent deficits (Dahmani and Bohbot, 2020), suggesting that recovery may depend on the nature of the skill and the type of assistance provided.
Several promising avenues emerge from our limitations. Longitudinal studies extending beyond four weeks could reveal whether skill degradation occurs or whether reliance on AI continues or whether recovery occurs after AI assistance withdrawal. Cross-cultural validation would establish whether dependency effects vary across different educational systems, media landscapes, or cultural approaches to authority and expertise. Comparative studies examining different AI pedagogical approaches — particularly Socratic questioning methods that prompt user reasoning rather than providing direct answers — could identify design principles that enhance rather than undermine skill development.
No AI condition. While our design effectively measures the impact of AI dialogue on misinformation discernment, a no-AI control condition performing the same longitudinal task would provide additional insights into whether humans become worse at discerning misinformation due to skill degradation or reliance on AI. Future work should examine this comparison to find the reason.
@@ 12. Conclusion
This study reveals a concerning paradox in AI-assisted misinformation detection: while persuasive AI systems produce immediate accuracy gains (+21%), they simultaneously undermine long-term independent judgment (15.3% by week 4 compared to week 0). Our longitudinal analysis demonstrates that current approaches prioritize belief correction over skill development, creating dependency rather than durable discernment capabilities. The findings challenge assumptions about AI’s educational potential and highlight the need for systems designed to enhance rather than replace human reasoning. As AI becomes increasingly sophisticated, ensuring these tools build critical thinking skills rather than cognitive dependency becomes essential for maintaining public resilience to misinformation.
13. Data Availability
All data, code, and materials from this study are available on GitHub (https://github.com/mitmedialab/newsapp).
Acknowledgements.
The authors would like to thank the Media Lab Consortium, MIT Tata Center Technology and Design Fellowship, and Google PhD Fellowship in Human–Computer Interaction for supporting the work. The authors would also like to thank Twishmay Shankar and Chanakya Ekbote for early feedback and discussions.References
- AI suggestions homogenize writing toward western styles and diminish cultural nuances. In Proceedings of the 2025 CHI Conference on Human Factors in Computing Systems, pp. 1–21. Cited by: §10.2, §2.4.
- The psychology of rumor.. Cited by: §2.1.
- Open deep search: democratizing search with open-source reasoning agents. arXiv preprint arXiv:2503.20201. Cited by: §3.1.
- Fake news, fast and slow: deliberation reduces belief in false (but not true) news headlines.. Journal of experimental psychology: general 149 (8), pp. 1608. Cited by: §2.1.
- Human vs. machine: a comparative study on the detection of ai-generated content. ACM Transactions on Asian and Low-Resource Language Information Processing 24 (2), pp. 1–26. Cited by: §1.
- Transforming qualitative information: thematic analysis and code development. Sage. Cited by: §9.
- To trust or to think: cognitive forcing functions can reduce overreliance on ai in ai-assisted decision-making. Proceedings of the ACM on Human-computer Interaction 5 (CSCW1), pp. 1–21. Cited by: §1, §2.4, §2.4.
- Endoscopist deskilling risk after exposure to artificial intelligence in colonoscopy: a multicentre, observational study. The Lancet Gastroenterology & Hepatology 10 (10), pp. 896–903. Cited by: §10.2.
- Advancing knowledge together: integrating large language model-based conversational ai in small group collaborative learning. In Extended Abstracts of the CHI Conference on Human Factors in Computing Systems, CHI EA ’24, New York, NY, USA. External Links: ISBN 9798400703317, Link, Document Cited by: §1.
- Can a humorous conversational agent enhance learning experience and outcomes?. In Proceedings of the 2021 CHI Conference on Human Factors in Computing Systems, CHI ’21, New York, NY, USA. External Links: ISBN 9781450380966, Link, Document Cited by: §1.
- Using conversational agents to support learning by teaching. External Links: 1909.13443, Link Cited by: §1.
- Did the roll-out of community notes reduce engagement with misinformation on x/twitter?. Proc. ACM Hum.-Comput. Interact. 8 (CSCW2). External Links: Link, Document Cited by: §1.
- Did the roll-out of community notes reduce engagement with misinformation on x/twitter?. Proceedings of the ACM on human-computer interaction 8 (CSCW2), pp. 1–52. Cited by: §1, §2.2.
- [14] Just the facts: how dialogues with ai reduce conspiracy beliefs.[preprint](2025). Cited by: §2.3.
- Durably reducing conspiracy beliefs through dialogues with ai. Science 385 (6714), pp. eadq1814. Cited by: §1, §1, §10.1, §2.3, §3.1, §5.
- Research design: qualitative, quantitative, and mixed methods approaches. Sage publications. Cited by: §9.
- Habitual use of gps negatively impacts spatial memory during self-guided navigation. Scientific reports 10 (1), pp. 6310. Cited by: §11.
- Deceptive explanations by large language models lead people to change their beliefs about misinformation more often than honest explanations. In Proceedings of the 2025 CHI Conference on Human Factors in Computing Systems, pp. 1–31. Cited by: §2.1.
- Wearable reasoner: towards enhanced human rationality through a wearable device with an explainable ai assistant. In Proceedings of the Augmented Humans International Conference, pp. 1–12. Cited by: §2.2, §2.4.
- Don’t just tell me, ask me: ai systems that intelligently frame explanations as questions improve human logical discernment accuracy over causal ai explanations. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems, pp. 1–13. Cited by: §1, §10.2, §10.4, §2.4.
- Algorithm aversion: people erroneously avoid algorithms after seeing them err.. Journal of experimental psychology: General 144 (1), pp. 114. Cited by: §2.4.
- Reminders and repetition of misinformation: helping or hindering its retraction?. Journal of applied research in memory and cognition 6 (2), pp. 185–192. Cited by: §2.2.
- The psychological drivers of misinformation belief and its resistance to correction. Nature Reviews Psychology 1 (1), pp. 13–29. Cited by: §2.1.
- Explicit warnings reduce but do not eliminate the continued influence of misinformation. Memory & cognition 38 (8), pp. 1087–1100. Cited by: §1, §2.1, §2.1, §2.2.
- Deepfakes and media integrity: navigating the new reality of synthetic content. Journal of Media and Interdisciplinary Studies 3 (9). Cited by: §2.1.
- Knowledge does not protect against illusory truth.. Journal of experimental psychology: general 144 (5), pp. 993. Cited by: §2.1.
- AI tools in society: impacts on cognitive offloading and the future of critical thinking. Societies 15 (1), pp. 6. Cited by: §1.
- How algorithms create and prevent fake news. Springer. Cited by: §2.1.
- Feeds of distrust: investigating how ai-powered news chatbots shape user trust and perceptions. ACM Transactions on Interactive Intelligent Systems. Cited by: §2.3.
- AI-driven mediation strategies for audience depolarisation in online debates. In Proceedings of the 2024 CHI Conference on Human Factors in Computing Systems, pp. 1–18. Cited by: §2.3.
- Consequences of cognitive offloading: boosting performance but diminishing memory. Quarterly Journal of Experimental Psychology 74 (9), pp. 1477–1496. Cited by: §11.
- Deepfake detection by human crowds, machines, and machine-informed crowds. Proceedings of the National Academy of Sciences 119 (1), pp. e2110013119. Cited by: §2.1.
- A survey on llm-as-a-judge. arXiv preprint arXiv:2411.15594. Cited by: §7.
- People are more susceptible to misinformation with realistic ai-synthesized images that provide strong evidence to headlines. Harvard Kennedy School Misinformation Review. Cited by: §2.1.
- The effects of repetition frequency on the illusory truth effect. Cognitive research: principles and implications 6 (1), pp. 38. Cited by: §2.1.
- Learning about informal fallacies and the detection of fake news: an experimental intervention. PLoS One 18 (3), pp. e0283238. Cited by: §2.1.
- MiRAGeNews: multimodal realistic ai-generated news detection. arXiv preprint arXiv:2410.09045. Cited by: Figure 4, §3.2.
- Does evaluating belief prior to its retraction influence the efficacy of later corrections?. Memory 28 (5), pp. 617–631. Cited by: §2.1, §2.2.
- Detection of ai-generated image content in news and journalism. In 2024 15th International Conference on Computing Communication and Networking Technologies (ICCCNT), pp. 1–6. Cited by: §2.1.
- Sources of the continued influence effect: when misinformation in memory affects later inferences. Journal of Experimental Psychology: Learning, Memory, and Cognition 20 (6), pp. 1420–1436. External Links: Document Cited by: §2.1, §2.1.
- Sources of the continued influence effect: when misinformation in memory affects later inferences.. Journal of experimental psychology: Learning, memory, and cognition 20 (6), pp. 1420. Cited by: §1.
- Thinking, fast and slow. Farrar, Straus and Giroux, New York, NY. Cited by: §2.1.
- Book review: speech and language processing by daniel jurafsky and james h. martin. Computational Linguistics 35 (3). Cited by: §8.
- Your brain on chatgpt: accumulation of cognitive debt when using an ai assistant for essay writing task. arXiv preprint arXiv:2506.08872 4. Cited by: §1, §10.2, §2.4.
- In generative ai we trust: can chatbots effectively verify political information?. Journal of Computational Social Science 8 (1), pp. 15. Cited by: §1.
- The impact of generative ai on critical thinking: self-reported reductions in cognitive effort and confidence effects from a survey of knowledge workers. In Proceedings of the 2025 CHI conference on human factors in computing systems, pp. 1–22. Cited by: §2.4.
- Misinformation and its correction: continued influence and successful debiasing. Psychological Science in the Public Interest 13 (3), pp. 106–131. External Links: Document Cited by: §2.1, §2.2.
- Algorithm appreciation: people prefer algorithmic to human judgment. Organizational Behavior and Human Decision Processes 151, pp. 90–103. Cited by: §2.4.
- Foundations of statistical natural language processing. MIT press. Cited by: §8.
- LLM-take: theme-aware keyword extraction using large language models. In 2023 IEEE International Conference on Big Data (BigData), pp. 4318–4324. Cited by: §8.
- Webgpt: browser-assisted question-answering with human feedback. arXiv preprint arXiv:2112.09332. Cited by: §3.1.
- The role of text representation in students’ ability to identify fallacious arguments. The Quarterly Journal of Experimental Psychology Section A 56 (5), pp. 849–864. Cited by: §2.1.
- Nonprobative photographs (or words) inflate truthiness. Psychonomic Bulletin & Review 19 (5), pp. 969–974. Cited by: §2.1.
- Humans and automation: use, misuse, disuse, abuse. Human factors 39 (2), pp. 230–253. Cited by: §2.4.
- Prior exposure increases perceived accuracy of fake news.. Journal of experimental psychology: general 147 (12), pp. 1865. Cited by: §2.1.
- Shifting attention to accuracy can reduce misinformation online. Nature 592 (7855), pp. 590–595. Cited by: §2.2.
- Fighting covid-19 misinformation on social media: experimental evidence for a scalable accuracy-nudge intervention. In Proceedings of the National Academy of Sciences, Vol. 117, pp. 19858–19866. External Links: Document Cited by: §2.2.
- Fighting misinformation on social media using crowdsourced judgments of news source quality. Proceedings of the National Academy of Sciences 116 (7), pp. 2521–2526. External Links: Document, Link, https://www.pnas.org/doi/pdf/10.1073/pnas.1806781116 Cited by: §2.1.
- Lazy, not biased: susceptibility to partisan fake news is better explained by lack of reasoning than by motivated reasoning. Cognition 188, pp. 39–50. Cited by: §1.
- The psychology of fake news. Trends in cognitive sciences 25 (5), pp. 388–402. Cited by: §2.2.
- Accuracy prompts are a replicable and generalizable approach for reducing the spread of misinformation. Nature communications 13 (1), pp. 2333. Cited by: §2.2.
- The elaboration likelihood model of persuasion. In Advances in experimental social psychology, Vol. 19, pp. 123–205. Cited by: §2.1.
- Manipulating and measuring model interpretability. In Proceedings of the 2021 CHI conference on human factors in computing systems, pp. 1–52. Cited by: §2.4.
- The continued influence of implied and explicitly stated misinformation in news reports.. Journal of experimental psychology: learning, memory, and cognition 42 (1), pp. 62. Cited by: §2.1, §2.2.
- Empowering learners: chatbot-mediated ’learning-by-teaching’. In Extended Abstracts of the CHI Conference on Human Factors in Computing Systems, CHI EA ’24, New York, NY, USA. External Links: ISBN 9798400703317, Link, Document Cited by: §1.
- Listening to misinformation while driving: cognitive load and the effectiveness of (repeated) corrections.. Journal of Applied Research in Memory and Cognition 12 (3), pp. 325. Cited by: §2.1, §2.2.
- De-fake: detection and attribution of fake images generated by text-to-image generation models. In Proceedings of the 2023 ACM SIGSAC conference on computer and communications security, pp. 3418–3432. Cited by: §1.
- People over trust ai-generated medical responses and view them to be as valid as doctors, despite low accuracy. arXiv e-prints, pp. arXiv–2408. Cited by: §1.
- Pictures and repeated exposure increase perceived accuracy of news headlines. Applied Cognitive Psychology 34 (5), pp. 1061–1071. Cited by: §2.1.
- Actively open-minded thinking and its measurement. Journal of Intelligence 11 (2), pp. 27. Cited by: §2.1.
- Intuition, reason, and metacognition. Cognitive psychology 63 (3), pp. 107–140. Cited by: §2.1.
- Judgment under uncertainty: heuristics and biases: biases in judgments reveal some heuristics of thinking under uncertainty.. science 185 (4157), pp. 1124–1131. Cited by: §2.1.
- Deepfakes and disinformation: exploring the impact of synthetic political video on deception, uncertainty, and trust in news. Social media+ society 6 (1), pp. 2056305120903408. Cited by: §2.1.
- Inoculating the public against misinformation about climate change. Global challenges 1 (2), pp. 1600008. Cited by: §2.2.
- The spread of true and false news online. Science 359 (6380), pp. 1146–1151. External Links: Document, Link, https://www.science.org/doi/pdf/10.1126/science.aap9559 Cited by: §2.1.
- Fact-checking: a meta-analysis of what works and for whom. Political Communication 37 (3), pp. 350–375. External Links: Document, Link, https://doi.org/10.1080/10584609.2019.1668894 Cited by: §1.
- A meta-analytic examination of the continued influence of misinformation in the face of correction: how powerful is it, why does it happen, and how to stop it?. Communication research 47 (2), pp. 155–177. Cited by: §2.1, §2.2.
- CNN-generated images are surprisingly easy to spot… for now. In Proceedings of the IEEE/CVF conference on computer vision and pattern recognition, pp. 8695–8704. Cited by: §1.
- Dire for diffusion-generated image detection. In Proceedings of the IEEE/CVF International Conference on Computer Vision, pp. 22445–22455. Cited by: §1.
- An overview of online fake news: characterization, detection, and discussion. Information Processing & Management 57 (2), pp. 102025. Cited by: §2.1.
- Judging llm-as-a-judge with mt-bench and chatbot arena. Advances in neural information processing systems 36, pp. 46595–46623. Cited by: §7.
@@ D.2. Conversation Characteristics
The conversations typically followed a pattern where participants would state their initial belief about a news item’s authenticity, the AI would probe their reasoning, provide evidence or counterevidence, and engage in iterative dialogue aimed at improving the participant’s detection accuracy. The relatively consistent conversation length across weeks indicates stable engagement patterns throughout the longitudinal study. Some of the examples of these conversations are present in Figure 13, 14 and 15.